OpenStudy (anonymous):

Friends!!!!! Help me with this log- sove for x : Ix-10I log(base2) x-3 = 2(x-10)

OpenStudy (anonymous):

\[\left| x-10 \right| \log_{2}(x-3) = 2(x-10)\]

OpenStudy (anonymous):

your questions i must say at one sight can give a person heart attackkkkkk wait i am solving

OpenStudy (anonymous):

are the two answers 7 and 13/4

OpenStudy (anonymous):

13/4 and 10

OpenStudy (anonymous):

Have you got 2 cases for solving mod??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

7 is not satisfying the equation

OpenStudy (anonymous):

see i solve the question you tell me where i am wrong

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

case 1: when x>10,mod(x-10)=x-10

OpenStudy (anonymous):

so x-10 cancels on both sides we get,,,, log(base 2)(x-3)=2 x-3=4 x=7

OpenStudy (anonymous):

but x > 10 therefore, it is discarded.

OpenStudy (anonymous):

ya i forgot

OpenStudy (anonymous):

in second case, when x<10, we have mod(x-10)=-(x-10) -(x-10)log(base 2)(x-3)=2(x-10) -log(base 2)(x-3)=2 log(base 2)1-log(base 2)(x-3)=2 log(base 2)(1/x-3)=2 1/x-3=4 x-3=1/4 x=13/4

OpenStudy (anonymous):

log(base 2)1-log(base 2)(x-3)=2 how???

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i understod

OpenStudy (anonymous):

because log(base 2)1=0 as log 1=0

OpenStudy (anonymous):

i am not able to find the error so that x=10

OpenStudy (anonymous):

second case is fine right?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

ok thanks!!! help me with one more question ia am posting.

OpenStudy (anonymous):

ok